OpenStudy (anonymous):

daniel ate 1/3 apple today 1/2 tomorrow and when he looked in bag had 2 apples left. how many apples did he begin with?

OpenStudy (mathteacher1729):

Do you mean: day 1 -- he ate one third of one apple day 2 -- he ate one half of one apple day 2 -- (after eating) he had exactly 2 apples in his bag?

OpenStudy (anonymous):

x-1/3x-1/2x=2 where x is the number of apples \[x(1-1/3-1/2)=2\] \[x(1/6)=2\] \[x=12\] however, your question is really not clear, your language is ambiguous. but if this is clear enough for you as an answer then fan me plz.

OpenStudy (anonymous):

there was 12 aples in the bad in beggining